# Worth Their Weight: Randomized and Regularized Block Kaczmarz Algorithms without Preprocessing

### v1

@@ F.1 Experiments with random matrices
The experiments in Sections 3.3 and 4.2 differ only in how the matrix $A$ is generated. For the left panel in Figure 1, we generate the matrix $A$ by setting each entry independently as $A_{ij}\sim\mathcal{N}(0,1)$. For the right panel, we construct $A=GU$ where $G\in\mathbb{R}^{m\times n}$ has independent standard normal entries and and $U$ has $\sigma_{i}=1/i^{2}$ and random orthonormal singular vectors. The condition number of the matrix $A$ is $\kappa(A)\approx 1$ in the left panel and $\kappa(A)\approx 10^{4}$ in the right panel.
The examples in Figure 2 are constructed as discretizations of underlying continuous problems, which is both a realistic scenario and serves to ensure that $A$ contains many nearly singular blocks. In both cases, we first generate an $n\times n$ matrix $C$. For the left panel we set $C=I$, whereas for the right panel we set $C$ with singular values $\sigma_{i}=1/i^{2}$ and random orthonormal singular vectors. Once $C$ is constructed, we define a set of $n$ functions $f_{1},\ldots,f_{n}$ by
$f_{j}(s)=\sum_{\ell=1}^{n}C_{j\ell}T_{\ell}(s),$ (101)
@@
$v_{i}=\left(-1+2\frac{i-1}{m-1}\right).$ (102)
@@
F.1.1 Effect of Regularization Parameter $\lambda$
In Figure 4, we investigate the effect of the parameter $\lambda$ on the performance of ReBlocK. The problem is the same as in the right panel of Figure 2, in which $A$ contains many ill-conditioned blocks $A_{S}$ and exhibits rapid singular value decay. It is observed that choosing $\lambda=1$ approximately reproduces the results of mSGD, while $\lambda=1e-3,1e-6,1e-9$ converges much faster, with $\lambda=1e-9$ beginning to show signs of instability. The results suggest that the performance of ReBlocK is only moderately sensitive to the choice of $\lambda$, since the parameter must be varied by multiple orders of magnitude to drastically change the performance.
Figure 4: Performance of TA-ReBlocK on the problem from the right panel of Figure 2 with various values of $\lambda$. The reported quantity is the suboptimality of the approximate solution $x$ for the unweighted problem (1), namely the value of $\epsilon$ for which $\left\lVert Ax-b\right\rVert=(1+\epsilon)\left\lVert r\right\rVert$. The vertical dotted line indicates the burn-in period of $T_{b}=300$, before which results are shown for individual iterates.
@@ F.2 Natural Gradient Experiments
$q(s)=\frac{1}{\sqrt{d}}\sum_{\ell=1}^{d}c_{\ell}T_{\ell}(s)$ (106)
for $d=15$ and each $c_{\ell}$ chosen randomly from a standard normal distribution. The resulting function is pictured in Figure 5.
@@
The neural network model is a simple periodic ResNet (He et al., 2016) with 5 layers. For input $s\in\mathbb{R}$, the network outputs $f(s)\in\mathbb{R}$ given by
$y_{1}=\tanh(W_{1}\sin(2\pi s)+b_{1}),$ (107)
$y_{i}=y_{i-1}+\tanh(W_{i}y_{i-1}+b_{i});\;i=2,3,4,$ (108)
$f(s)=W_{5}y_{4}+b_{5}.$ (109)
@@
The intermediate layers have dimensions $y_{i}\in\mathbb{R}^{50}$ and the weight matrices $W_{i}$ and bias vectors $b_{i}$ have the appropriate dimensions to match. The weights are initialized using a Lecun normal initialization and the biases are all initialized to zero. The parameters are collected into a single vector $\theta\in\mathbb{R}^{7801}$ for convenience, leading to the neural network function $f_{\theta}(r)$.
@@
The loss function is defined as in (26) and the network is trained using subsampled natural gradient descent, as described for example in (Ren & Goldfarb, 2019), with a batch size of $N_{b}=200$, a Tikhonov regularization of $\lambda=0.01$, and a step size of $\eta=0.5$. This corresponds to the parameter update
$\theta_{t+1}=\theta_{t}-\eta J_{S}^{\top}(J_{S}J_{S}^{\top}+N_{b}\lambda I)^{-1}[f_{\theta}-f]_{S},$ (110)
@@
where $S$ represents the set of $N_{b}$ sample points. The setting of $N_{b}=200$ is intended to represent something close to a full batch training regime, which is only practical since we are studying a very small network on a very compact, low-dimensional domain. The resulting training curve is presented in Figure 6.
@@
Figure 6: Training curve for the neural network function regression example. The three vertical lines indicate the training snapshots that are used to generate the least-squares problems studied in Figure 3.
@@
From this training run, three snapshots are taken and used to generate the least-squares problems for Figure 3, following Equation 31. The first snapshot is from the “pre-descent” phase before the loss begins to decrease, the second is from the “descent” phase during which the loss decreases rapidly, and the third is from the “post-descent” phase when the loss has stopped decreasing significantly. The snapshots are indicated by the vertical dotted lines in Figure 6. The algorithms TA-mSGD, TA-RBK, and TA-ReBlocK are then tested on the resulting problems wth an initial guess of $x_{0}=0$ and a batch size of $k=30$. The batch size of $k=30$ is meant to represent the realistic scenario when each iteration uses too few samples to thoroughly represent the target function. Furthermore, the continuous problems are treated directly by uniformly sampling $k=30$ points from the domain $[0,1]$ at each iteration and calculating the network outputs and gradients at these points.
@@
The step size for minibatch SGD is constant within each run and has been tuned independently for each example, specifically to be as large as possible without introducing any signs of instability, up to a factor of $2$. The value of $\lambda$ for ReBlocK has been set to $\lambda=0.001$ and is not optimized on a per-example basis. The reported quantity is the relative residual $\tilde{r}=\left\lVert Jx-[f_{\theta}-f]\right\rVert/\left\lVert f_{\theta}-f\right\rVert$, which measures how well the function-space update direction $Jx$ agrees with the function-space loss gradient $f_{\theta}-f$. This quantity is estimated accurately by sampling $m=2\cdot 10^{4}$ points from $[0,1]$. Furthermore, since it is expensive to calculate this residual, the quantity is only calculated and reported at approximately $10^{3}$ evenly spaced iterations for each run of each algorithm.
F.2.1 Iteration Cost of Each Algorithm
In Figure 7, we report the number of iterations per second for mSGD, RBK, and ReBlocK for the experiments of Section 5. We find that RBK iterations are approximately five times slower than ReBlocK iterations and six times slower than mSGD iterations for these particular problems. Interestingly, there is only a slight difference in speed between the ReBlocK and mSGD iterations. The experiments are conducted in single precision on an A100 GPU to simulate a deep learning setting.
Figure 7: Iteration cost of each algorithm for the natural gradient experiments of Section 5.
F.2.2 Singular Value Analysis
In Figure 8, we analyze the singular values of the Jacobian operator $J$ for each experiment in Figure 3. We first approximate $J$ by a finite matrix $A$ by sampling $m=2\cdot 10^{4}$ points from the domain $[0,1]$, then use SciPy’s svds function with the ARPACK solver to approximate the top $200$ singular values of $A$. We find that in every case the top singular values decay exponentially up to some threshold, after which the tail decays more slowly. The initial exponential decay helps to explain the faster convergence rate of ReBlocK relative to mSGD.
Figure 8: Top $200$ singular values of the input matrix $A$ for each natural gradient-based least-squares problem. The left, middle, and right panels correspond to the same panels in Figure 3, and the vertical dotted line represents the batch size of $k=30$ used in the least-squares solvers for Figure 3.

### v2

@@ F.1 Experiments with random matrices
The experiments in Sections 3.3 and 4.2 differ only in how the matrix $A$ is generated. Thus, we first describe the matrix construction for each case separately, then describe the rest of the set-up in a unified manner.
Beginning with the experiments of Section 3.3, for the left panel of Fig. 1 we generate the matrix $A$ by setting each entry independently as $A_{ij}\sim\mathcal{N}(0,1)$. For the right panel, we construct $A=GU$ where $G\in\mathbb{R}^{m\times n}$ has independent standard normal entries and and $U$ has $\sigma_{i}=1/i^{2}$ and random orthonormal singular vectors. The condition number of the matrix $A$ is $\kappa(A)\approx 1$ in the left panel and $\kappa(A)\approx 10^{4}$ in the right panel.
Regarding Section 4.2, the examples in Fig. 3 are constructed as discretizations of underlying continuous problems, which is both a realistic scenario and serves to ensure that $A$ contains many nearly singular blocks. In both cases, we first generate an $n\times n$ matrix $C$. For the left panel we set $C=I$, whereas for the right panel we set $C$ with singular values $\sigma_{i}=1/i^{2}$ and random orthonormal singular vectors. Once $C$ is constructed, we define a set of $n$ functions $f_{1},\ldots,f_{n}$ by
(101) $f_{j}(s)=\sum_{\ell=1}^{n}C_{j\ell}T_{\ell}(s),$
@@
(102) $v_{i}=\left(-1+2\frac{i-1}{m-1}\right).$
@@ F.2 Natural Gradient Experiments
(106) $q(s)=\frac{1}{\sqrt{d}}\sum_{\ell=1}^{d}c_{\ell}T_{\ell}(s)$
for $d=30$ and each $c_{\ell}$ chosen randomly from a standard normal distribution. The resulting function is pictured in Fig. 9.
@@
The neural network model is a simple periodic ResNet [21] with 5 layers. For input $s\in\mathbb{R}$, the network outputs $f(s)\in\mathbb{R}$ given by
(107) $y_{1}=\tanh(W_{1}\sin(2\pi s)+b_{1}),$
(108) $y_{i}=y_{i-1}+\tanh(W_{i}y_{i-1}+b_{i});\;i=2,3,4,$
(109) $f(s)=W_{5}y_{4}+b_{5}.$
@@
The intermediate layers have dimensions $y_{i}\in\mathbb{R}^{500}$ and the weight matrices $W_{i}$ and bias vectors $b_{i}$ have the appropriate dimensions to match. The weights are initialized using a Lecun normal initialization and the biases are all initialized to zero. The parameters are collected into a single vector $\theta\in\mathbb{R}^{753,001}$ for convenience, leading to the neural network function $f_{\theta}(r)$.
@@
The loss function is defined as in (27) and the network is trained using subsampled natural gradient descent, as described for example in [38], with a batch size of $N_{b}=500$, a Tikhonov regularization of $\lambda=0.01$, and a step size of $\eta=0.5$. This corresponds to the parameter update
(110) $\theta_{t+1}=\theta_{t}-\eta J_{S}^{\top}(J_{S}J_{S}^{\top}+N_{b}\lambda I)^{-1}[f_{\theta}-f]_{S},$
@@
where $S$ represents the set of $N_{b}$ sample points. The setting of $N_{b}=500$ is intended to represent something close to a full batch training regime, which is only practical since our function regression problem lives in a very compact, one-dimensional domain. The resulting training curve is presented in Fig. 10.
@@
Figure 10: Training curve for the neural network function regression example. The three vertical lines indicate the training snapshots that are used to generate the least-squares problems studied in Fig. 5.
@@
From this training run, three snapshots are taken and used to generate the least-squares problems for Fig. 5, following Eq. 32. The first snapshot is from the “pre-descent” phase before the loss begins to decrease, the second is from the “descent” phase during which the loss decreases rapidly, and the third is from the “post-descent” phase when the decay rate of the loss has slowed substantially. The snapshots are indicated by the vertical dotted lines in Fig. 10 and form the basis of the experiments of Figs. 5, 6, 7, and 8. The batch size of $k=50$ used in Fig. 5 is meant to represent the realistic scenario when each iteration uses too few samples to thoroughly represent the target function. Furthermore, the continuous problems are treated directly by uniformly sampling $k$ points from the domain $[0,1]$ at each iteration and calculating the network outputs and gradients at these points.
@@
The initial guess is $x_{0}=0$ in every case and the step size for minibatch SGD is constant within each run and has been tuned independently for each example, specifically to be as large as possible without introducing any signs of instability, up to a factor of $2$. The value of $\lambda$ for ReBlocK has been set to $\lambda=0.001$ unless otherwise specified, and is not optimized on a per-example basis. The reported quantity is the relative residual $\tilde{r}=\left\lVert Jx-[f_{\theta}-f]\right\rVert/\left\lVert f_{\theta}-f\right\rVert$, which measures how well the function-space update direction $Jx$ agrees with the function-space loss gradient $f_{\theta}-f$. This quantity is estimated at each iteration using the available sample points.
Disclaimer